The vapor pressure, P, of a certain liquid was measured at two temperatures If you were going to graphically determine the enthalpy of vaporizaton, AHvap, for this liquid, what points would you plot? The data is shown in this table Number Number T(K) P (kPa) 225 3.75 675 8,62 To avoid rounding errors, use three significant figures in the x values and four significant figures in the y values. point 1 Number Number NOTE: Keep the pressure units in kPa. point 2:n Determine the rise, run, and slope of the line formed by these points rise run slope Number Number Number What is the enthalpy of vaporization of this liquid? Number J/mol vap
Answer #1

here graph is drawn between ln P and 1/ T

lnp = y -axis

1/T = x - axis

point 1:   x = 4.44 x 10^-3   ,   y = 1.322

point 2 : x = 1.48 x 10^-3   , y = 2.154

rise =   y2-y1 = 0.832

run = x2 - x1    = - 2.96 x 10^-3

slope = rise / run = - 0.832 / 2.96 x 10^-3 = - 281.1

H / R = slope

\DeltaH = 281.1 x 8.314 = 2337 J / mol